Suppose a consumer consuming two commodities X and Y has the following utility function X0.4 Y0.6. If price of good X and Y are 2 and 3 respectively and income constraints birr 50.

A/ Find the quantities of X and Y which Maximize utility

B/ Find the MRSxy


1
Expert's answer
2021-08-20T08:48:13-0400

Solution:

A.). Quantities of X and Y which maximize utility:

Utility maximizing condition is where "\frac{MUx} {MUy} = \frac{Px} {Py}"

U(x,y) = 10X0.4Y0.6

Determine MUx and MUy:

MUx = "\frac{\partial U} {\partial X}" = 4X0.4-1Y0.6 = 4X– 0.6Y0.6

 

MUy = "\frac{\partial U} {\partial Y}" = 6X0.4 Y0.6-1 = 6X0.4Y-0.4

 

Set "\frac{MUx} {MUy} = \frac{Px} {Py}" :

Px = 2

Py = 3

"\frac{4X^{-0.6}Y^{0.6} } {6X^{0.4}Y^{-0.4}} = \frac{2} {3}"


Simplify:

"\frac{2Y} {3X} = \frac{2} {3}"


Y = X


Derive the budget constraint:

Budget constraint: I = PxX + PyY

50 = 2X + 3Y

Substitute the X value in the budget constraint to derive Y:

50 = 2X + 3X

50 = 5X

X = 10

Since Y = X, then Y = 10

The quantities of X and Y that maximizes utility (Ux,y) are = (10,10)


B.). Find the MRSxy:

MRSxy = MUx/MUy = "\frac{4X^{-0.6}Y^{0.6} } {6X^{0.4}Y^{-0.4}} = \frac{2Y} {3X} = \frac{2(10)} {3(10)} = \frac{20} {30} = \frac{2} {3}" =20/30 = 2/3

MRSxy = "\frac{2} {3}" or 0.67